OpenStudy (anonymous):

Find all solutions of the system of equations. x−16y^2 =0 4y-x^2= 0

OpenStudy (experimentx):

one solution will be 0,0

OpenStudy (experimentx):

substitute the value of x from first equation on the second equation x−16y^2 =0 => x = 16y^2 4y-x^2= 0 4y-(16y^2)^2= 0 4y(1 - 64y^3) = 0 y^3 = 1/64 y = 1/4 putting this value on first equation, x = 16*(1/4)^2 = 1 so your two real solutions are 0,0 and 1,1/4 theoretically it should have 4 solutions, so other solution are imaginary.

OpenStudy (experimentx):

complete solution is here http://www.wolframalpha.com/input/?i=x%E2%88%9216y%5E2+%3D0%2C+4y-x%5E2%3D+0

OpenStudy (anonymous):

Awesome. Thank you!

OpenStudy (experimentx):

You are welcome!!!